OpenStudy (anonymous):

simplifying triginometry identities: 1/cot x-1/2sinx secx

geerky42 (geerky42):

\(\dfrac{1}{\cot x}-\dfrac{1}{2\sin x \sec x}\) ?

OpenStudy (anonymous):